Naranjo Company designs industrial prototypes for outside companies. Budgeted overhead for the year was $260,000, and budgeted direct labor hours were 20,000. The average wage rate for direct labor is expected to be $25 per hour. During June, Naranjo Company worked on four jobs. Data relating to these four jobs follow:

Job 39 Job 40 Job 41 Job 42
Beginning balance $23,700 $34,600 $17,000 $0
Materials requisitioned 18,900 21,400 8,350 12,000
Direct labor cost 10,000 18,500 3,000 2,900

Overhead is assigned as a percentage of direct labor cost. During June, Jobs 39 and 40 were completed; Job 39 was sold at 130 percent of cost. (Naranjo had originally developed Job 40 to order for a customer; however, that customer was near bankruptcy and the chance of Naranjo being paid was growing dimmer. Naranjo decided to hold Job 40 in inventory while the customer worked out its financial difficulties. Job 40 is the only job in Finished Goods Inventory.) Jobs 41 and 42 remain unfinished at the end of the month.

Required:

1. Calculate the overhead rate based on direct labor cost.

% of direct labor cost

2. Set up a simple job-order cost sheet for all jobs in process during June. If an amount is zero, enter "0".

Naranjo Company
Job-Order Cost Sheets
Job 39 Job 40 Job 41 Job 42
Balance, June 1 $ $ $ $
Total $ $ $ $

3. What if the expected direct labor rate at the beginning of the year was $20 instead of $25? What would the overhead rate be?

New budgeted direct labor cost = $

New overhead rate = % of direct labor cost

How would the cost of the jobs be affected?
Answer #1

PART-1) Solution: 0.52

Working: Budgeted direct labor cost = average wage rate for direct labor x budgeted direct labor hours 25*20,000

Overhead rate = budgeted overhead costs ÷ budgeted direct labor cost = 260,000 / 500,000 = 0.52

PART-2)

Job 39

Job 40

Job 41

Job 42

Beginning balance

$23,700

$34,600

$17,000

$0

Material requisitioned

$18,900

$21,400

$8,350

$12,000

Direct labor cost

$10,000

$18,500

$3,000

$2,900

Applied Overhead

$5,200

$9,620

$1,560

$1,508

Total cost

$57,800

$84,120

$29,910

$16,408

Part-3) What would the overhead rate be? $1

Working: =260000/(20*20000)
Answer #2

3. What if the expected direct labor rate at the beginning of the year was $20 instead of $25? What would the overhead rate be?

New budgeted direct labor cost = $

New overhead rate = % of direct labor cost

How would the cost of the jobs be affected?

Applied overhead would increase 

Workout: 

New budgeted direct labor cost= 20*200,000

New overhead rate= 260,000/(20*200,000)
source: Notes
answered by: Stephanie